OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) HELP INEQUAILTY

OpenStudy (anonymous):

get rid of the parentheses first using the distributive property

OpenStudy (anonymous):

@satellite73 okay so then we would have -2x + 6 ≥ 5 – 3x?

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

left side is correct, you would get \(-2x+6\) but the right side should be \(5-x-3=2-x\)

OpenStudy (saifoo.khan):

@SchoolFreak : same question, again and again. :( http://openstudy.com/users/schoolfreak#/updates/50263871e4b07eea166b2889

OpenStudy (anonymous):

that is, \(5 – (x + 3)=5-x-3=2-x\) by the distributive law

OpenStudy (saifoo.khan):

similar*

OpenStudy (anonymous):

so start with \[ -2x + 6 ≥ 2-x\] then add \(2x\) to both sides, and subtract \(2\) from both sides that you will be done

OpenStudy (anonymous):

6 + 4 ≥ 4 - x and then what im confused for this part ?? @satellite73

OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) -2x+6>= 5-x-3 6-2x>=2-x 6-2>=-x+2x x<=4

OpenStudy (anonymous):

(look here careflly) 6-2x>=2-x 6-2>=-x+2x x<=4

OpenStudy (anonymous):

how did you do it explain in words??

OpenStudy (anonymous):

-2x+6>= 5-x-3 Did u understand up to here

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now add -2x to both sides: -2x+6>= 5-x-3 -2x+6+2x>=2-x+2x

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

so what do u get?

OpenStudy (anonymous):

-2x+6+2x>=2-x+2x after this

OpenStudy (anonymous):

6-2x>=2-x

OpenStudy (anonymous):

add +2x on both sides

OpenStudy (anonymous):

6-2>=-x+2x

OpenStudy (anonymous):

????

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

now, 4>=x

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

and thats the answer

OpenStudy (anonymous):

@SchoolFreak got it?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

how did you get 4 >= x

OpenStudy (anonymous):

6-2>=-x+2x (from there)

OpenStudy (anonymous):

6-2=4 (right?0